OpenStudy (anonymous):

One number is 7 more than twice another. If their difference is 22, what is the larger number?

OpenStudy (anonymous):

\[x=7+2y\] \[x -y=22\] therefore \[22+y=7+2y\] \[y=15\] then sub into one of the equations and solve for x

OpenStudy (anonymous):

btw let x= larger number and y= lower number

OpenStudy (anonymous):

29 37 43 here are the answer choices

OpenStudy (anonymous):

if u subbed it in its 37